# TomoGraphView: 3D Medical Image Classification with Omnidirectional Slice Representations and Graph Neural Networks

### v1

@@ 4.8 Comparative Analysis - 33D Pretrained Models
In the following, we compare our proposed TomoGraphView framework with large-scale pretrained 3D models. All compared models are convolution-based, as vision transformer architectures for 3D images are highly memory-intensive due to the substantially increased token length (i.e., number of patches) relative to 2D images. Moreover, relative to convolutional methods, vision transformers typically require much larger datasets for effective pretraining, while medical imaging data generally remains comparatively limited.
@@
The 3D models included in our comparison with TomoGraphView correspond to the top-performing approaches for downstream classification identified in the studies by aerts2025foundation [aerts2025foundation] and wald2024openmind [wald2024openmind]. In particular, we benchmark our framework against FMCIB [pai2024foundation], Vista3D [he2025vista3d], Models Genesis [zhou2019models], and two pretrained models introduced by wald2024openmind [wald2024openmind] based on VoCo [wu2024voco] and SwinUNETR [tang2022self]. A detailed description of the models is provided in C.
@@
The remainder of this section is structured as follows. In Section 4.8.1, we first evaluate the fixed embeddings obtained from large-scale pretrained 3D models across all datasets, without applying any task-specific finetuning to assess their representational capacity learned during pretraining. Subsequently, in Section 4.8.2, we perform dataset-specific finetuning, which typically improves downstream performance by allowing the pretrained representations to better adapt to the characteristics of each target task. Since all large-scale pretrained 3D models were trained in a self-supervised manner, we attach a classification head to the encoder to enable downstream classification. Consistent with our previous experiments, we fix the number of trainable parameters to 100100k to ensure a fair comparison. Notably, prediction heads previously discussed, such as GNN, transformer, or LSTM-based architectures, cannot be integrated with the 33D pretrained models, as they require slice-based inputs rather than full 3D volume embeddings.
4.8.1 Frozen Backbones
@@
We begin by evaluating the five large-scale pretrained 3D models across all six oncological downstream tasks and compare their performance with our proposed TomoGraphView framework. Results are depicted in the left radar plot of Figure 8. Detailed quantitative results across methods, datasets, and performance metrics are presented in Appendix Table 6.
@@
Across all tasks, the proposed TomoGraphView framework achieves an average AUROC of $0.8282$, outperforming the on average second-best approach, FMCIB [pai2024foundation], which achieves an AUROC of $0.7170$. This demonstrates a consistent advantage of TomoGraphView over all competing models. Looking more closely, on a per-task level, TomoGraphView also yields the highest AUROC for each dataset: breast tumors ($0.7214$ vs. $0.6207$, SwinUNETR), brain tumors ($0.9639$ vs. $0.8985$, FMCIB), head–neck tumors ($0.6999$ vs. $0.5910$, SwinUNETR), kidney tumors ($0.7750$ vs. $0.7466$, SwinUNETR), liver tumors ($0.9664$ vs. $0.8711$, FMCIB), and soft-tissue tumors ($0.8426$ vs. $0.6926$, SwinUNETR). Although the performance gains are less pronounced for kidney tumors ($0.7750$ vs. $0.7466$) and brain tumors ($0.9639$ vs. $0.8985$), TomoGraphView consistently outperforms all competing methods by a clear margin across the remaining breast, head-neck, liver, and soft-tissue tumor datasets. This observation, as exemplified by the AUROC results, can be further confirmed by additional evaluation metrics, such as balanced accuracy, F1-Score, and Matthews correlation coefficient, underscoring the great potential of TomoGraphView. Detailed results can be found in the Appendix Table 6.
4.8.2 33D Model Fine-tuning Performance
Building on the frozen backbone results presented in the previous subsection, where FMCIB [pai2024foundation] emerged as the best-performing large-scale pretrained 3D model on average across all datasets in terms of AUROC, balanced accuracy, F1-score, and Matthews correlation coefficient, we now evaluate its finetuning performance, as it allows the pretrained representations to better adapt to the specific characteristics of each downstream task. Strictly following wald2024openmind [wald2024openmind], we finetune FMCIB for 200 epochs in total, using a learning rate of $1\times 10^{-4}$, a cosine annealing scheduler, a gradual warm-up of $100$ epochs applied to both the encoder and the classification head, and a batch size of 256. Additionally, we explore a batch size of $16$, which is consistent with previous experiments. The corresponding results are illustrated in the right radar plot of Figure 8, with detailed quantitative results provided in Appendix Table 7, which additionally shows decreased warm-up durations of $20$ and $50$ epochs to allow for more update steps under higher learning rates.
@@ 5 Discussion
In this work, we introduce TomoGraphView, a novel framework designed for 33D medical image classification that preserves the spatial structure of volumetric medical imaging data while enabling compatibility with 22D feature encoders through both canonical and non-canonical 33D volume slicing and graph-based feature aggregation. We comprehensively evaluate our approach across six diverse oncology datasets covering multiple anatomical regions (e.g., brain, breast, head-neck, kidney, or liver), various imaging modalities (CT and multiple MRI sequences), and various classification tasks primarily concerning tumor characterization. Building on our two key contributions, omnidirectional volume slicing and spherical mesh-based feature aggregation, our experimental evaluations show that TomoGraphView achieves state-of-the-art results across downstream classification tasks, and even outperforms large-scale pretrained 33D models. The following section provides a detailed discussion of the results and findings obtained throughout this work.
5.1 Why Omnidirectional Volume Slicing Matters
As the first key contribution of this work, we introduce our novel omnidirectional volumetric slicing strategy, which extends beyond axial, coronal, and sagittal orientations, thereby enabling a richer characterization of volumetric structures that are not naturally aligned with canonical axes. The value of rich characterization through both meaningful and representative volume slicing to describe target 33D structures is supported by our experimental evaluations in Section 3.1, which demonstrate that incorporating additional planes, transitioning from 22D axial to 2.52.5D slicing, consistently enhances downstream performance. On average, AUROC performance across all evaluation datasets improves by $4.71\%$ (from $0.7335$ to $0.7681$) when comparing 2.52.5D slicing with its corresponding $3$-view 22D-axial+ configuration. This trend also persists for a larger number of views: with 24 views, AUROC increases by $2.39\%$ (from $0.7701$ to $0.7885$). However, critically, further gains are achieved with omnidirectional volume slicing, which improves AUROC by $3.41\%$ (from $0.7885$ to $0.8154$) compared to 2.5D2.5D+ slicing, and by $5.88\%$ (from $0.7701$ to $0.8154$) relative to pure axial slicing.
@@ 5.1 Why Omnidirectional Volume Slicing Matters
Notably, the benefit of omnidirectional volume slicing generalizes beyond MLPs across various prediction head backbones, including LSTM, and transformer-based architectures, as outlined in Section 4.7. Interestingly, although our omnidirectional slicing strategy cannot be directly converted into a naturally meaningful slice ordering, as compared to consecutive axial slicing, LSTMs trained on omnidirectional slices outperform those trained on consecutive axial slices, with a notable AUROC increase of $5.60\%$ (from $0.7446$ to $0.7863$). Similarly, transformer-based models show an AUROC improvement of $5.83\%$ (from $0.7746$ to $0.8198$).
Overall, the superior results obtained throughout this work demonstrate that omnidirectional slicing provides a richer representation of volumetric structures, yielding consistent and architecture-agnostic performance enhancements. These findings highlight the versatility and added value of our proposed slicing strategy as a compelling alternative to traditional approaches.
5.2 Graph-based Feature Aggregation Excels Baselines
@@
Beyond the choice of representing a volume as a set of 22D slices, where omnidirectional volume slicing emerged as the primary choice throughout our experiments, the way information from cross-sectional images is aggregated is of central importance. To this end, we present our second core contribution in this work, GNN-based feature aggregation, which relies on a spherical mesh-based graph topology to explicitly model the spatial relationships of encoded 22D slices extracted from 33D volumes. Our spherical graph design is derived from the omnidirectional slicing strategy, enabling the integration of information across a set of in-plane and out-of-plane perspectives while explicitly encoding spatial relationships within the graph topology. Our experimental evaluations in Section 4.4 and Section 4.7 underscore the advantage of our graph-based feature aggregation method across datasets and backbones.
@@
While MLPs do not offer any built-in capability to model relationships among slices, LSTMs treat consecutive slices as temporal sequences, thereby preserving their spatial ordering within a volume. However, such an ordering is only naturally defined along a single plane of consecutive slices, restricting the representation to one direction and potentially overlooking complementary information from other orientations. Nevertheless, although omnidirectional volume slicing improves LSTM downstream performance, there is no natural ordering of these slices, potentially hindering further improvements.
@@
We attribute the superior performance of our GNN-based approach to its stronger inductive bias. Specifically, the node feature aggregation and node connectivity are predefined, whereas for transformers, such as MST-based aggregation, spatial relationships and global dependencies must be learned through their self-attention mechanism, introducing additional complexity. Moreover, transformers typically perform best in large-scale data settings, which are uncommon in medical imaging, limiting their practical suitability. We argue that simplifying the learning task by the topological constraints introduced in our GNN-based approach leads to improved overall average performance while handling limited data more efficiently.
5.3 Dimensionality-Quantity Tradeoff in 3D Classification
@@
Current approaches to 33D medical image classification face a fundamental trade-off between data dimensionality and data quantity. Although 22D slice-based methods achieve strong performance, they inherently lose anatomical context and spatial coherence when 33D volumes are decomposed into individual slices. Conversely, models that directly process 33D data preserve spatial structure but are constrained by the limited availability of pretraining data. Both dimensionality and data scarcity substantially impact a model’s representational capacity for volumetric analysis, a crucial aspect for reliable downstream classification.
@@
While foundation models have advanced rapidly in domains such as histopathology [filiot2024phikon, chen2024towards, koch2024dinobloom], where large-scale 22D data and sophisticated pretraining strategies drive strong transferability, native 33D foundation models remain comparatively underdeveloped [van2025foundation]. Our findings in Section 4.8 support this observation.
@@
We begin by comparing TomoGraphView with five large-scale pretrained 33D models across six oncology datasets, using frozen encoder backbones. Evaluating raw embeddings enables a resource-efficient and faithful assessment of the representational quality learned during pretraining. Ideally, a robust pretrained model should produce embeddings that capture rich image semantics and correlate well with downstream performance. As TomoGraphView also relies solely on frozen DINOv2 features, this comparison remains fair. Across all datasets, the FMCIB model [pai2024foundation] performs best among 33D approaches but consistently trails TomoGraphView, with relative improvements of 15.5% in AUROC and 16.8% in balanced accuracy (see Table 6 in the Appendix).
@@
Interestingly, despite being pretrained exclusively on body CT data (DeepLesion [yan2018deeplesion]), FMCIB even outperforms MRI-pretrained models on brain tumor MRI classification and matches the performance of the MRI-pretrained SwinUNETR [wald2024openmind]. We attribute this cross-modality generalization to the model’s ability to learn universal imaging features such as textures, edges, and spatial hierarchies.
@@
We further evaluate fine-tuning performance to assess how well FMCIB adapts pretrained representations to specific downstream tasks. Following the fine-tuning setup of wald2024openmind [wald2024openmind], we observe improvements for brain, soft-tissue, and head–neck tumors, but no gain for breast, kidney, or liver datasets. This highlights that the effectiveness of fine-tuning is highly dataset-dependent and that no single configuration generalizes universally without extensive hyperparameter optimization. Nevertheless, overall fine-tuning performance still falls short of TomoGraphView. In this study, we focused exclusively on finetuning large-scale pretrained 33D models, as transfer learning has consistently demonstrated superior performance over training from scratch, particularly in medical image classification tasks where labeled data is often scarce. Existing studies show that leveraging prior knowledge from large-scale datasets enables faster convergence and improved generalization to domain-specific medical data [tajbakhsh2016convolutional].
@@
In summary, the results from Sections 4.8.1 and 4.8.2 highlight the added value of combining 22D foundation models with advanced volume-slicing and feature-aggregation strategies. We therefore position TomoGraphView as a key step toward bridging the gap until fully native 33D foundation models become available for medical image analysis.

### v2

@@ 4.8 Comparative Analysis with 33D Pretrained Models
In the following, we compare our proposed TomoGraphView framework at $24$ views with large-scale pretrained 3D models. All baseline models are convolution-based, as 3D vision transformers are substantially more memory-demanding due to the much larger token sequences (i.e., number of patches) compared to 2D inputs. In addition, vision transformers typically require significantly larger datasets to pretrain effectively, whereas medical imaging datasets are often relatively limited in size.
@@
The 3D models used in our comparison with TomoGraphView represent the top-performing downstream classification approaches identified by aerts2025foundation [aerts2025foundation] and wald2024openmind [wald2024openmind]. Specifically, we benchmark our framework against FMCIB [pai2024foundation], Vista3D [he2025vista3d], Models Genesis [zhou2019models], as well as two pretrained models from wald2024openmind [wald2024openmind] based on VoCo [wu2024voco] and SwinUNETR [tang2022self]. A detailed description of these models is provided in C.
@@
The remainder of this section is structured as follows. In Section 4.8.1, we first evaluate the fixed embeddings obtained from the selected large-scale pretrained 3D models across all datasets, without applying any task-specific finetuning to assess their representational capacity learned during pretraining. Subsequently, in Section 4.8.2, we perform dataset-specific finetuning, which typically improves downstream performance by allowing the pretrained representations to better adapt to the characteristics of each target task. Since all large-scale pretrained 3D models were trained in a self-supervised manner, we attach a classification head to the encoder to enable downstream classification. Consistent with our previous experiments, we fix the number of trainable parameters to 100100k to ensure a fair comparison. Notably, prediction heads previously discussed, such as GNN, transformer, or LSTM-based architectures, cannot be integrated with the 33D pretrained models, as they require slice-based inputs rather than full 3D volume embeddings.
4.8.1 TomoGraphView vs. Frozen 3D Backbones
@@
We first evaluate the five large-scale pretrained 3D models across all six oncological downstream tasks and compare their performance with our proposed TomoGraphView framework. The results are summarized in the left radar plot of Figure 8. Comprehensive quantitative results for all methods, datasets, and performance metrics are provided in Appendix Table 5.
@@
Across all tasks, our proposed TomoGraphView framework achieves an average AUROC of $0.8372$, outperforming the on average second-best approach, FMCIB [pai2024foundation], which achieves an AUROC of $0.7170$. This demonstrates a consistent advantage of TomoGraphView over all competing models. Looking more closely, on a per-task level, TomoGraphView also yields the highest AUROC for each dataset: breast tumors ($0.7084$ vs. $0.6207$, SwinUNETR), brain tumors ($0.9655$ vs. $0.8985$, FMCIB), head–neck tumors ($0.7527$ vs. $0.5910$, SwinUNETR), kidney tumors ($0.7893$ vs. $0.7466$, SwinUNETR), liver tumors ($0.9634$ vs. $0.8711$, FMCIB), and soft-tissue tumors ($0.8441$ vs. $0.6926$, SwinUNETR). Although the performance gains are less pronounced for kidney tumors ($0.7893$ vs. $0.7466$) and brain tumors ($0.9655$ vs. $0.8985$), TomoGraphView consistently outperforms all competing methods by a clear margin across the remaining breast, head-neck, liver, and soft-tissue tumor datasets. This observation, as exemplified by the AUROC results, can be further confirmed by additional evaluation metrics, such as balanced accuracy, F1-Score, and Matthews correlation coefficient, underscoring the great potential of TomoGraphView. Detailed results can be found in the Appendix Table 5.
4.8.2 TomoGraphView vs. Finetuned 3D Backbones
Building on the frozen backbone results presented in the previous subsection, where, on average, the FMCIB [pai2024foundation] approach emerged as the best-performing large-scale pretrained 3D model across all datasets and different performance metrics, we now evaluate its finetuning performance, as it allows the pretrained representations to better adapt to the specific characteristics of each downstream task. Strictly following the scheme introduced by wald2024openmind [wald2024openmind], we finetune FMCIB for $200$ epochs in total, using a learning rate of $1\times 10^{-4}$, a cosine annealing scheduler, a gradual warm-up of $100$ epochs applied to both the encoder and the classification head, and a batch size of $256$. Additionally, we explore a batch size of $16$, which is consistent with previous experiments. The corresponding results are illustrated in the right radar plot of Figure 8, with detailed quantitative results provided in Appendix Table 6, which additionally shows decreased warm-up durations of $20$ and $50$ epochs to allow for more update steps under higher learning rates.
@@ 5 Discussion
In this work, we introduce TomoGraphView, a novel framework designed for 33D medical image classification that preserves the spatial structure of volumetric medical imaging data while enabling compatibility with powerful 22D vision foundation models. We comprehensively evaluate our approach across six diverse oncology datasets covering multiple anatomical regions (e.g., brain, breast, head-neck, kidney, or liver), various imaging modalities (CT and multiple MRI sequences), and various classification tasks primarily concerning tumor characterization. Building on our two key contributions, omnidirectional volume slicing and spherical mesh-based feature aggregation, our experimental evaluations show that TomoGraphView achieves state-of-the-art results across downstream classification tasks, and even outperforms large-scale pretrained 33D models. The following section provides a detailed discussion of the results and findings obtained throughout this work.
5.1 Advancing Volume Slicing using Omnidirectional Views
Generally, our experimental results shown in Table 1 suggest that more expressive tumor characterization, achieved by using richer and more representative volume slices from complementary directions, improves the description of the underlying 3D target structures. Specifically, expanding from purely axial 22D slices to a 2.52.5D multi-plane strategy consistently yields better downstream performance. As a result, the average AUROC performance across all evaluation datasets improves by $4.71\%$ (from $0.7335$ to $0.7681$) when comparing 2.52.5D slicing with its corresponding 22D-axial+ configuration (i.e., with three axial slices). This trend also persists for a larger number of views: with 24 views, AUROC increases by $2.39\%$ (from $0.7701$ to $0.7885$).
@@ 5.1 Advancing Volume Slicing using Omnidirectional Views
However, as the first key contribution of this work, we presented a novel omnidirectional volumetric slicing strategy that extends beyond the three conventional orientations axial, coronal, and sagittal and enables a richer characterization of volumetric structures that are not inherently aligned with these canonical axes. As a consequence, further gains are achieved with omnidirectional volume slicing, which improves AUROC by $3.41\%$ (from $0.7885$ to $0.8154$) compared to 2.5D2.5D+ slicing, and by $5.88\%$ (from $0.7701$ to $0.8154$) relative to pure 22D-axial slicing. It is worth highlighting, that the benefit of omnidirectional volume slicing generalizes beyond MLP prediction heads across various backbones, including LSTM-, transformer-, and GNN-based architectures, as outlined in Section 4.7. Notably, although our omnidirectional slicing strategy cannot be directly converted into a naturally meaningful slice ordering, as compared to consecutive axial slicing, LSTMs trained on omnidirectional slices outperform those trained on consecutive axial slices, with a notable AUROC increase of $5.60\%$ (from $0.7446$ to $0.7863$). Similarly, transformer-based models show an AUROC improvement of $5.83\%$ (from $0.7746$ to $0.8198$).
While our omnidirectional volume slicing strategy consistently improves the average performance across all experiments in Table 1, it does not always yield the single best score for every individual dataset, although it generally remains very close. We attribute these differences as reflections of intrinsic dataset-specific characteristics. In particular, as all datasets offer their highest in-plane resolution when extracting axial slices (i.e., finer spacing in the coronal and sagittal directions), we examine the alignment of target structures along the z-axis. This axis exhibits the greatest anisotropy across all datasets and therefore governs the amount of low-resolution lesion information captured when slices are taken from predominantly coronal or sagittal orientations. To quantify anatomical z-axis alignment, we apply principal component analysis (PCA) to the 3D segmentation masks to identify each structure’s major axis. The resulting alignment score is computed as the absolute cosine similarity between this major axis and the anatomical z-axis, yielding a value between 0 and 1, corresponding to no alignment and perfect alignment, respectively. Z-axis alignment scores across all datasets are visible in the Appendix Table 7. Overall, we observe the following trend: For datasets with high z-axis alignment such as head-neck (0.92) and soft-tissue tumors (0.93), omnidirectional volume slicing strongly improves downstream performance, owing to its inherent slicing plane variability. However, datasets which show a lower z-axis alignment such as breast (0.40) and liver (0.10) tumors naturally benefit from several high-resolution axial slices as offered by e.g., 2.52.5D+ volume slicing.
Taken together, the strong results throughout this work show that omnidirectional slicing captures volumetric structures more effectively than conventional slicing strategies, leading to consistent, architecture-independent performance gains. These outcomes underscore the flexibility and added benefit of our approach as a compelling alternative to traditional methods such as 22D-axial or 2.52.5D slicing. We argue that omnidirectional slicing is highly relevant for the medical image analysis community working with slice-based pipelines, as it provides a simple yet elegant way to represent 3D targets in a spatially coherent manner. To facilitate adoption, we also provide a user-friendly Python library for omnidirectional volume slicing at https://pypi.org/project/OmniSlicer.
5.2 Spatially Coherent Graph-Based Feature Aggregation
@@
Beyond the choice of representing a volume as a set of 22D slices, the way information from cross-sectional images is aggregated is of central importance. To this end, we presented our second core contribution of this work: GNN-based feature aggregation, which relies on a spherical mesh-based graph topology to explicitly model the spatial relationships of encoded 22D slices extracted from 33D volumes. Our spherical graph design is derived from the omnidirectional slicing strategy, enabling the integration of information across a set of in-plane and out-of-plane perspectives while explicitly encoding their spatial relationships within the graph topology. Our experimental evaluations in Section 4.4 and Section 4.7 underscore the advantage of our graph-based feature aggregation method across datasets.
@@
While MLPs do not offer any built-in capability to model relationships among slices, LSTMs treat consecutive slices as temporal sequences, thereby preserving their spatial ordering within a volume. However, such an ordering is only naturally defined along a single plane of consecutive slices, restricting the representation to one direction and potentially overlooking complementary information from other orientations. Nevertheless, although omnidirectional volume slicing improves LSTM downstream performance, there is no natural ordering of these slices, potentially hindering further gains.
@@
We attribute the superior performance of our GNN-based approach to its stronger inductive bias. Specifically, the node feature aggregation and node connectivity are predefined, whereas for transformers, such as MST-based aggregation, spatial relationships and global dependencies must be learned through their self-attention mechanism, introducing additional complexity. Moreover, transformers typically perform best in large-scale data settings, which are uncommon in medical imaging, limiting their practical application. We argue that simplifying the learning task by the topological constraints introduced in our GNN-based approach leads to improved overall average performance while handling limited data more efficiently.
Furthermore, we argue that the performance improvements also lie in the GNN’s ability to explicitly integrate both local and global information by propagating slice-level features across nodes through edge connections defined by the graph structure (see Section 3.3). This mechanism is not available e.g., to the MLP, which instead processes all slice information through fully connected layers without explicit relational modeling. Furthermore, given that the employed graph topology approximates a three-dimensional geometric structure (i.e., a sphere), we argue that a graph-based representation is better suited to preserve the inherent spatial relationships among slices compared to dense connections in an MLP. Consequently, the graph structure provides a more faithful representation of the 33D volume when represented through 22D slices and their corresponding relative position.
5.3 Balancing Data Dimensionality and Data Scarcity
@@
Current approaches to 33D medical image classification face a fundamental trade-off between data dimensionality (2D slices vs. 3D volumes) and data quantity (i.e., sample size). Although 22D slice-based methods achieve strong performance, they inherently lose anatomical context and spatial coherence when 33D volumes are decomposed into individual slices. Conversely, models that directly process 33D data preserve spatial structure but are constrained by the limited availability of pretraining data. Both, data dimensionality and data scarcity substantially impact a model’s representational capacity for volumetric analysis, two crucial aspects for reliable downstream classification.
@@
While foundation models have advanced rapidly in domains such as histopathology [filiot2024phikon, chen2024towards, koch2024dinobloom], where large-scale 22D data and sophisticated pretraining strategies drive strong transferability, radiology data-based 33D foundation models remain comparatively underdeveloped [van2025foundation]. Our findings in Section 4.8 support this observation.
@@
We began by comparing TomoGraphView with five large-scale pretrained 33D models across six oncology datasets, using frozen encoder backbones. Evaluating raw embeddings enables a resource-efficient and faithful assessment of the representational quality learned during pretraining. Ideally, a robust pretrained model should produce embeddings that capture rich image semantics and correlate well with downstream performance. As TomoGraphView also relies solely on frozen DINOv2 features, this comparison remains fair. Across all datasets, the FMCIB model [pai2024foundation] performs best among 33D approaches but consistently trails TomoGraphView, with relative improvements of 15.5% in AUROC and 16.8% in balanced accuracy (see Appendix Table 5).
@@
Interestingly, despite FMCIB pai2024foundation being pretrained exclusively on body CT data (DeepLesion [yan2018deeplesion]), FMCIB even outperforms MRI-pretrained models on brain tumor MRI classification and matches the performance of the MRI-pretrained SwinUNETR [wald2024openmind]. We attribute this cross-modality generalization to the model’s ability to learn universal imaging features such as textures, edges, and spatial hierarchies.
@@
We further evaluated fine-tuning performance to assess how well FMCIB adapts pretrained representations to specific downstream tasks. Closely following the fine-tuning setup of wald2024openmind [wald2024openmind], we observe improvements for brain, soft-tissue, and head–neck tumors, but no gain for breast, kidney, or liver datasets. This highlights that the effectiveness of fine-tuning is highly dataset-dependent and that no single configuration generalizes universally without extensive hyperparameter optimization. Nevertheless, overall fine-tuning performance still falls short of TomoGraphView. In this study, we focused exclusively on finetuning large-scale pretrained 33D models, as transfer learning has consistently demonstrated superior performance over training from scratch, particularly in medical image classification tasks where labeled data is often scarce. Existing studies show that leveraging prior knowledge from large-scale datasets enables faster convergence and improved generalization to domain-specific medical data [tajbakhsh2016convolutional].
@@
In summary, the results from Sections 4.8.1 and 4.8.2 highlight the added value of combining 22D foundation models with advanced volume-slicing and feature-aggregation strategies. We therefore position TomoGraphView as a key component toward bridging the gap until fully native 33D foundation models become available for medical image analysis.